OpenStudy (bbb911):

Veterinarian Jasmine Cato has 2/9 of a quart of medication. She wishes to prescribe this medication for 3 cats in her pet hospital. If she divides the medication evenly, how much will each cat receive?

OpenStudy (anonymous):

If you are dividing some quantity into n even parts, then the resulting quantities will be 1/n of that original quantity. You arrive at tha tquantity by multiplying the original quantity by 1/n. In your case, you have a division by 3, so from that information, you should be able to get what "n" is and what 1/n is. Take that substituted 1/n and multiply that by the original quantity. That will give you the individual amount for each cat.

OpenStudy (anonymous):

Also,\[\frac{ a }{ b } \times \frac{ 1 }{ n } = \frac{ a }{ b \times n }\]

OpenStudy (bbb911):

i am kind of confused

OpenStudy (anonymous):

So, if you have "n" even parts and there are 3 cats, what is "n"?

OpenStudy (anonymous):

It's not a trick question. It's something rather clear, and the question is asked to you to get you to start. So, go ahead and come up with "n". Give it a try.

OpenStudy (bbb911):

nvmind i got all i did was divide 2/9 by 3 i got 4/15

OpenStudy (anonymous):

It is true that you want to divide 2/9 by 3, but the answer is not 4/15. If you want to use a method other than the one I outlined, that's fine, but something went wrong with the way you did it. You might want to show your work so I can help you see where you went wrong.

OpenStudy (bbb911):

yea i actually did do it right i typed in the answer said it was correct. this is only basic math there's no algerba yet. it just finally clicked to me

OpenStudy (anonymous):

If you don't like my first method of solving, here's another: Multiply the numerator and denominator by 3/3. Then take 1/3 of the resulting numerator. In other words, the original numerator. This amount to multiplying taking the original numerator for the numerator answer and original denominator times 3 for the denominator answer.

OpenStudy (bbb911):

wait a minute im so dumb right now i accidentally switched to a different question omg sorry bout tht i got the right answer for a different answer

OpenStudy (anonymous):

np. Now that you are putting question and answer together on the same problem now, to divide a fraction into "n" equal parts, you take that fraction's numerator and keep it the same. You multiply the denominator by "n" for the resulting denominator. That's a third way to look at the problem and perhaps the easiest. There, that's 3 ways to look at it. If you want, you take any of the methods, but you might want to stick with this last one if you are stuck with the others.

OpenStudy (bbb911):

yea the real question was this 4/5 and 3 thts why i got 4/15

OpenStudy (anonymous):

yes, that's correct. 1/3 of 4/5 is indeed 4/15. So, you got it right! Good job!

OpenStudy (bbb911):

yay!! thank you :) sorry for the trouble imma give you a medal for helping me to realize.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @bbb911 ! And your're welcome! Nice working with you!

OpenStudy (bbb911):

ur welcome :)